OpenStudy (anonymous):

FIND THE ROOTS AND MAKE A GRAPH A. y=x^2+5x+4 B. y=-2x^2+2

jhonyy9 (jhonyy9):

A y=(x+4)(x+1)

jhonyy9 (jhonyy9):

B y=2(x ^{2} +1) = 2((x+1)^{2}-2x)\]